# Dichography: Two-frame Ultrafast Imaging from a Single Diffraction Pattern

### v2

@@ Abstract
X-ray free-electron lasers can nowadays deliver pairs of ultrabright, ultrashort light pulses with controllable delays at different photon energies. When combined with diffraction experiments, the interaction of these pulses with a sample at different times can produce two snapshots of an evolving system, enabling spatial and temporal resolution of ultrafast dynamics in nanomatter at and beyond the terahertz time scale. However, light detectors are orders of magnitude slower. The diffraction signals from the two pulses overlap in the recorded data and cannot be separated using currently available analysis methods. Here, we address this challenge by introducing Dichography, a diffraction imaging technique that restores the two unique views of the sample from superimposed scattering signals. We apply Dichography to experimental diffraction patterns of isolated xenon-doped superfluid helium nanodroplets, imaged using time-delayed, two-color X-ray pulses at intensities near the detection limit. The reconstructed images provide evidence of the survival of the xenon structures up to 750 fs after the interaction with the first shot. The capabilities of Dichography are further explored by applying the method to data from a second experiment, in which we retrieve the images of two distinct silver nanoparticles intercepted by the same light pulse. Dichography enables a new class of experiments for spatially and temporally resolving ultrafast phenomena, making a significant step toward the original promise of X-ray free-electron lasers to capture ultrafast movies of nanomatter.
I Introduction
@@
Among diffraction-based techniques, single-particle Coherent Diffraction Imaging (CDI) stands out for its capability to retrieve the full electronic density of isolated nanostructures from a single diffraction image produced by an individual light pulse [6, 7]. Thus, since the pioneering era of XFELs [8], CDI has been an invaluable tool for capturing ultrafast snapshtos of fragile nanoscale objects in free-flight [9, 10, 11, 12, 13] and their light-driven dynamics [14, 15, 16, 17, 18].
@@
Thanks to recent technical developments, XFELs can deliver pairs of light pulses in the XUV or X-ray spectrum [19, 1, 20, 21, 22, 23, 24]. Both pulses retain the coherence, brightness, and ultrashort nature of conventional XFEL pulses; their wavelength can be independently tuned, and their arrival time at the interaction region can be controlled with high precision. This novel capability, known as two-color mode, unlocks a whole new class of dynamical investigations, where the XFEL delivers both the pump pulse, which triggers the dynamics in a pristine sample, and the probe pulse, which probes the evolved state. When the two-color capability of XFELs is combined with the high spatial resolution of diffraction imaging, the sample’s structure can be accessed at two different points in time. The femtosecond precision at which the time delay between the two light pulses can be controlled paves the way for two-frame movies of structural dynamics reaching the terahertz realm, and potentially well beyond.
@@
Significant effort had been devoted to overcoming this technical limitation by physically separating the two scattering images. First successful realizations achieved this by placing spectral filters [29, 30] or by directing the pulses onto two different detectors via optical devices [31]. Nevertheless, these solutions come with significant trade-offs, e.g., strongly reduced signal brightness, restricted range of viable photon energies, reduced control on the time delay, or partial loss of scattering information. Therefore, developing new methods to fully exploit two-color pulses across various diffraction-based experimental approaches remains a top priority for the scientific community.
@@
In this work, we introduce Dichography (from dichos, meaning “in two”), an imaging method that enables the reconstruction of two sample images from superimposed diffraction signals. Dichography, applicable under a variety of experimental conditions, unlocks the full imaging potential of two-color diffraction patterns by disentangling the scattering contributions in the recorded data, to reconstruct the two, distinct, views of the sample as “seen” by the individual pulses.
@@
We perform dichographic imaging for two different experimental conditions. In one case, a pair of light pulses of different wavelengths delivered by the European XFEL intercepts superfluid helium nanodroplets doped with xenon. Despite the low photon yield in the diffraction signal – for such a highly photon-demanding approach – successful reconstructions can be achieved for a subset of optimal diffraction patterns. The retrieved frames are two snapshots of the same xenon-doped helium droplet, delayed in time by $750\text{\,}\mathrm{fs}$. Similar structures can be identified in the two frames, indicating that extended damages to the sample due to the interaction with the first XFEL pulse only happen at longer time scales.
@@
In a second experimental case, pairs of silver nanoparticles are simultaneously present in the focus when single light pulses delivered by the SwissFEL reach the interaction region. When the two samples have enough spatial separation, the interference of their scattered fields is not resolved by the detector. In this case, the properties of the imaging problem, i.e. the mathematical connection between the sample and the acquired data, become analogous to the two-color situation. Therefore, Dichography can be employed to restore the images of the two particles, demonstrating its flexibility and performance under different conditions.
@@
The imaging quality achieved by Dichography is strictly connected to the present understanding of the underlying mathematical and numerical challenges, as well as to the current performance of XFELs. With this work, we aim at stimulating and encouraging research efforts in both directions. In this regard, the possible applications of Dichography with the currently available experimental possibilities are discussed, along with concepts that could be viable in the near future.
@@ II The Dichographic problem
In conventional CDI, the image recorded by the detector, $I(\vec{q})$, is the square amplitude of the scattered field $\psi(\vec{q})$. The field carries the full information on the sample density $\rho(\vec{x})$. However, only its magnitude $\left|\psi(\vec{q})\right|=\sqrt{I(\vec{q})}$ is experimentally accessible, while the field’s phase is lost. Therefore, it is necessary to retrieve the phase to obtain the sample image $\rho(\vec{x})$. In CDI, the phase retrieval problem is numerically solved by specifically designed phase retrieval algorithms [32, 7].
The imaging problem in Dichography is significantly more demanding. Indeed, Dichography deals with situations where the acquired scattering data $I(q)$ is the incoherent sum of the intensities of two independent scattered fields $\psi^{A}(\vec{q})$ and $\psi^{B}(\vec{q})$, that is:
$I(\vec{q})=I^{A}(\vec{q})+I^{B}(\vec{q})=\left|\psi^{A}(\vec{q})\right|^{2}+\left|\psi^{B}(\vec{q})\right|^{2}\,.$ (1)
The two fields encode two different electronic densities $\rho^{A}$ and $\rho^{B}$, which correspond to the two independent images of a sample. Thus, the dichographic imaging problem consists in retrieving the phase of $\psi^{A}(\vec{q})$ and $\psi^{B}(\vec{q})$, as well as their relative amplitude contribution to the recorded signal $I(\vec{q})$.
Fig. 1 provides an intuitive comparison between the use case of Dichography and those of single-particle CDI and Holography, using a double-slit diffraction experiment as a reference. As previously mentioned, conventional CDI, illustrated in Fig. 1a, addresses the reconstruction of a single isolated sample, denoted by $\rho$. In contrast, Holography, depicted in Fig. 1b, involves two distinct samples, $\rho^{A}$ and $\rho^{B}$. This configuration necessitates the interference of the two resulting scattering fields at the detector, which fundamentally differs from the Dichography approach shown in Fig. 1c.
@@
The imaging algorithm at the core of Dichography, which disentangles the scattering contributions and retrieves the two frames, is a fundamental result of this work. Its structure is derived from that of conventional iterative phase retrieval algorithms for single-particle CDI, adapted to accommodate the specific characteristics of Dichography. In fact, it comprises of two iterative phase retrieval processes, each retrieving a frame. The two processes are solely entangled when the experimental diffraction data is constrained: the two estimated intensities are rescaled such that their sum corresponds to the recorded scattering image. A detailed description of the algorithm’s workflow and its connection to the CDI counterpart is deferred to Appendix A. We now turn to an investigation of the performance of Dichography in two different experimental situations.
III Results
III.1 Seeing double from a single shot
@@
Figure 2: Example of dichographic reconstruction from a single diffraction pattern produced by a “double-hit”. The experimental diffraction data from an individual light pulse delivered by SwissFEL is shown in a). The two particles density retrieved from a) via Dichography are shown in b) and c). The two reconstructions correspond to two silver nanocubes, with edge length of around $100\text{\,}\mathrm{nm}$. Once the two particles have been reconstructed, the corresponding single-particle diffraction patterns can be calculated, shown in d) for the reconstruction in b), and in e) for the one reported in c). The central parts of the disentangled diffraction images in d) and e) are magnified in f) and g), respectively. The white arrows point at the same region of the diffraction image. All diffraction data in a), d), e), f) and g) are plotted with logarithmic color scale.
@@
The first set of results concerns the analysis of diffraction patterns produced by two particles struck by the same XFEL shot, known as “double-hits”. Although they are of lower scientific relevance compared to two-color diffraction images, they serve as a valuable probe for investigating the capabilities, performance, and key features of Dichography.
@@
Here, pairs of silver nanoparticles, simultaneously present in the XFEL focus, are intercepted by a single light pulse delivered by SwissFEL [33]. When the particles are far enough apart, the interference term between the two scattered fields is not resolved by the detector. The recorded diffraction image is, in practice, the incoherent sum of the two intensities, suitable for imaging with Dichography. The two retrieved frames correspond to the electron density of the two different particles. Further experimental information are provided in Appendix D.1.
@@
The first dichographic imaging is shown in Fig. 2. The images of the two particles in Fig. 2b and 2c are the two frames retrieved from the experimental diffraction pattern reported in Fig. 2a, without prior information regarding the particles’ shape, size, or orientation. The reconstructions reveal that the experimental image is produced by two silver nanocubes of slightly different sizes and alignment. Due to the cubic structure of the two samples and their orientation, each cube produces strong streaks with four-fold symmetry in the diffraction image. A total of eight streaks can be indeed identified in the diffraction pattern in Fig. 2a.
@@
It is possible to disentangle the two contributions to the scattering, i.e. to calculate the corresponding single-particle diffraction images from the two reconstructed frames. Fig. 2d and Fig. 2e report the two disentangled scattering contributions from the first (Fig. 2b) and second (Fig. 2c) frames, respectively. There, we can appreciate the capability of Dichography to separate the two contributions to the scattering superimposed in 2a, by reproducing the four-fold symmetry of the two cubes with different orientations. This is further enhanced by a closer view in the central part of the separated patterns where the strongest contribution to the scattering signal is concentrated, provided in Fig. 2f and 2g.
@@
From these zoomed-in representations of the disentangled scattering signals, it is also possible to identify few regions where the intensities are not fully disentangled. For example, the white arrows in Fig. 2f and Fig. 2g indicate the same coordinates in the two patterns. In Fig. 2g, this area correspond to a high intensity streak in the direction of cube facets, as visible from the reconstruction in Fig. 2c. Despite the absence of any related structural feature in the other reconstructed frame in Fig. 2b, its disentangled pattern in Fig. 2f still presents traces of this strong scattering signal. This signal marginally affects the inner density profile in Fig. 2c, slightly distorted in the direction of the cube facets of the other frame. We refer to this “cross-talking” between the two frames – where features in one create slight artifacts in the other due to strongly differing scattering intensity in specific directions – as ghosting. On the other hand, the intensities mistakenly assigned to the first frame in Fig. 2f are about two orders of magnitude lower than those correctly retrieved in Fig. 2g. This is evident from the logarithmic color map on the right and underscores the ability of Dichography to separate the two signals, even when their intensities differ strongly.
@@
Further reconstruction attempts on diffraction patterns extracted from the same dataset of Fig. 2 are presented in Fig. 3. These examples are selected to highlight how Dichography performs on different sample configurations. For instance, Fig. 3a shows a reconstruction which reveals a pair of nanostructures with completely different geometries, a cube and a triangle, strongly underlining how this method neither requires any specific constraint in terms of sample’s structure, nor structural links between the two frames. The reconstruction in Fig. 3b has an additional challenging aspect. Here, the spatial extension of the two samples differs significantly, with the smaller one being around $100\text{\,}\mathrm{nm}$ in maximum size and the second reaching around $200\text{\,}\mathrm{nm}$. As a consequence, the smaller sample could “fit” into the larger one, potentially rendering the separation of the two frames of the reconstruction a significantly harder task. In both Fig. 3a and Fig. 3b, ghosting artifacts are barely visible in the form of line-shaped density fluctuations aligned with the hard edges of the reconstruction in the other frame.

### v3

# arXiv:2508.20153v3[physics.optics] 27 Feb 2026

@@ Abstract
We experimentally demonstrate that pairs of time-delayed ultrabright and ultrashort X-ray pulses of two different colors, delivered by modern X-ray Free Electron Lasers, can provide two time-delayed snapshots of a sample. We introduce Dichography, a method that algorithmically separates the diffraction signals overlapping on the detector and independently retrieves the two images of the specimen. We employ Dichography to reconstruct two views of individual xenon-doped helium nanodroplets with $20\text{\,}\mathrm{nm}$ spatial resolution. The consistency of structures observed in both images at delays up to $750\text{\,}\mathrm{fs}$ provides evidence that, under these illumination conditions, significant structural damage only occurs at longer timescales. We further validate the method by imaging pairs of silver nanoparticles intercepted by the same light pulse. Dichography enables a new class of experiments across physics, chemistry, and materials science, making a significant step toward the original promise of X-ray free-electron lasers to capture ultrafast movies of nanomatter.
@@
Among diffraction-based techniques, single-particle Coherent Diffraction Imaging (CDI) stands out for its capability to retrieve the full electronic density of isolated nanostructures from a single diffraction image produced by an individual light pulse [6, 7]. Thus, since the pioneering era of XFELs [8], CDI has been an invaluable tool for capturing ultrafast snapshots of fragile nanoscale objects in free-flight [9, 10, 11, 12, 13] and their light-driven dynamics [14, 15, 16, 17, 18].
@@
Thanks to recent technical developments, XFELs can deliver pairs of light pulses in the XUV or X-ray spectrum [19, 1, 20, 21, 22, 23, 24]. Both pulses retain the coherence, brightness, and ultrashort nature of conventional XFEL pulses; their wavelength can be independently tuned, and their arrival time at the interaction region can be controlled with high precision. This novel capability, known as two-color mode, unlocks a whole new class of dynamical investigations, where the XFEL delivers both the pump pulse, which triggers the dynamics in a pristine sample, and the probe pulse, which probes the evolved state. When the two-color capability of XFELs is combined with the high spatial resolution of diffraction imaging, the sample’s structure can be accessed at two different points in time. The femtosecond precision at which the time delay between the two light pulses can be controlled paves the way for two-frame movies of structural dynamics approaching the Petahertz realm, and potentially well beyond.
@@
Significant effort has been devoted to overcoming this technical limitation by physically separating the two scattering images. First successful realizations achieved this by placing spectral filters [29, 30] or by directing the pulses onto two different detectors via optical devices [31] or different regions of the detector through holographic approaches [32]. Nevertheless, these solutions come with significant trade-offs, e.g., strongly reduced signal brightness, restricted range of viable photon energies, reduced control on the time delay, partial loss of scattering information or limited spatial resolution. Therefore, developing new methods to fully exploit two-color pulses across various diffraction-based experimental approaches remains a top priority for the scientific community.
@@
In this work, we introduce Dichography (from dichos, meaning “in two”), an imaging method that enables the reconstruction of two sample images from superimposed diffraction signals without any prior knowledge of the samples shape and size. While relying on an established numerical and theoretical framework [33], Dichography extends the approach to real experimental conditions, i.e., samples with unknown structures, noisy diffraction images and missing information in the center of the patterns. Thereby, it unlocks the first experimental demonstration of single-particle two-color Coherent Diffraction Imaging at XFELs, disentangling the scattering contributions in the recorded data and reconstructing two distinct views of the sample as “seen” by the individual X-ray colors.
@@
Superfluid helium nanodroplets doped with xenon are intercepted by two-color X-ray pulses delivered by the European XFEL at photon energies of $1.0\text{\,}\mathrm{keV}$ and $1.2\text{\,}\mathrm{keV}$. Despite the low photon yield in the diffraction signal for such a highly photon-demanding approach, successful reconstructions can be achieved for a subset of optimal diffraction patterns. The retrieved frames are two snapshots of the same xenon-doped helium droplet, delayed in time by $50\text{\,}\mathrm{fs}$ and $750\text{\,}\mathrm{fs}$. The retrieved embedded xenon structures appear to be unchanged within the spatial resolution of the experiment, indicating that extensive structural damage of the helium-embedded nanostructures due to the interaction with the first XFEL pulse only occurs at longer time scales.
@@
For benchmarking the Dichography method, a further experimental case is also presented. Here, pairs of silver nanoparticles are simultaneously present in the focus when single light pulses delivered by the SwissFEL reach the interaction region. When the two samples are sufficiently separated in space, the interference of their scattered fields is not resolved by the detector. The properties of the imaging problem, i.e., the mathematical connection between the sample and the measured diffraction, are analogous to the two-color situation, i.e., an incoherent superposition of two diffraction patterns. Therefore, Dichography can be employed to restore the images of the two particles. While the analysis of these double-hit patterns is of limited scope regarding the physical implications, it serves as a crucial testbed, allowing us to show and investigate the performance of the method on experimental data for optimal diffraction conditions, not affected by the limited pulse energies of XFELs in two-color operation mode, and for diverse structural properties of the samples.
@@
The imaging quality and resolution achieved by Dichography is strictly connected to the present understanding of the underlying mathematical and numerical challenges, as well as to the current performance of XFELs. With this work, we also aim at stimulating and encouraging research efforts in both directions. In this regard, the possible applications of Dichography with the currently available experimental possibilities are discussed, along with concepts that could be viable in the near future.
@@ I The Dichographic problem
In conventional CDI, the image recorded by the detector, $I(\mathbf{q})$, is the square amplitude of the scattered field $\psi(\mathbf{q})$. The field carries the full information on the sample density $\rho(\mathbf{x})$. However, only its magnitude $\left|\psi(\mathbf{q})\right|=\sqrt{I(\mathbf{q})}$ is experimentally accessible, while the field’s phase is lost. Therefore, it is necessary to retrieve the phase to obtain the sample image $\rho(\mathbf{x})$. In two dimensions, the image $\rho(\mathbf{x})$ is proportional to the projection of the electronic density along the beam propagation direction [6]. In CDI, the phase retrieval problem is numerically solved by specifically designed phase retrieval algorithms [34, 7].
The imaging problem in Dichography is significantly more demanding. Indeed, Dichography deals with situations where the acquired scattering data $I(q)$ is the incoherent sum of the intensities of two independent scattered fields $\psi^{A}(\mathbf{q})$ and $\psi^{B}(\mathbf{q})$, that is:
$I(\mathbf{q})=I^{A}(\mathbf{q})+I^{B}(\mathbf{q})=\left|\psi^{A}(\mathbf{q})\right|^{2}+\left|\psi^{B}(\mathbf{q})\right|^{2}\,.$ (1)
The two fields encode two different electronic densities $\rho^{A}$ and $\rho^{B}$, which correspond to the two independent images of a sample. Thus, the dichographic imaging problem consists of retrieving the phase of $\psi^{A}(\mathbf{q})$ and $\psi^{B}(\mathbf{q})$, as well as their relative amplitude contribution to the recorded signal $I(\mathbf{q})$.
Fig. 1 provides an intuitive comparison between Dichography, single-particle CDI and Holography, using a double-slit diffraction experiment as a reference. As previously mentioned, conventional CDI, as illustrated in Fig. 1a, addresses the reconstruction of a single isolated sample, denoted by $\rho$. In contrast, Holography, depicted in Fig. 1b, involves two distinct samples, $\rho^{A}$ and $\rho^{B}$. This configuration necessitates the interference of the two resulting scattering fields at the detector, which fundamentally differs from the Dichography approach shown in Fig. 1c.
@@
The imaging algorithm at the core of Dichography, which disentangles the scattering contributions and retrieves the two frames, is a fundamental result of this work. The possibility of algorithmically disentangling scattering contributions incoherently adding up on the detector has already been shown in previous works [33], stimulated by the problem of imaging for arrangements beyond crystallography [35]. Dichography extends this concept to the reconstructions of isolated samples with unknown structures, thus effectively enabling two-color single-particle coherent diffraction imaging. The dichographic imaging problem is less constrained than conventional CDI and, thus, inherently harder. Still, in most practical cases, it is demonstrated to have a unique solution [33], as further discussed in Sec. V.2.
The algorithmic structure is derived from that of conventional iterative phase retrieval algorithms for single-particle CDI [36, 34, 37], adapted to accommodate the specific characteristics of Dichography, extending the scheme proposed in Ref. [33] to the retrieval of samples with unknown shapes.
A detailed description of the algorithm’s workflow and its connection to the CDI counterpart is deferred to Sec. V.1. We now turn to an investigation of the performance of Dichography in two different experimental situations.
II Results
II.1 Two-color Coherent Diffraction Movies of Xenon-doped Helium Nanodroplets
Figure 2: Reconstructions from two-color diffraction patterns acquired at the European XFEL, produced by superfluid helium nanodroplets doped with xenon. Two experimental diffraction patterns are reported in a) and d), on which the diffraction signals produced by two different photon energies, $1.0\text{\,}\mathrm{keV}$ and $1.2\text{\,}\mathrm{keV}$, are superimposed. The intensity is encoded as reported by the color bar at the bottom. The intensity values are translated into numbers of $1.2\text{\,}\mathrm{keV}$ photons per pixel. The inset plot is a zoomed-in region of the diffraction image, to better appreciate the photon statistics, and thus the statistical noise, that affects the experimental data. The two colors are delayed by $50\text{\,}\mathrm{fs}$ for the pattern in a) and $750\text{\,}\mathrm{fs}$ for the pattern in d). The two reconstructed frames, corresponding to the two time-delayed projections of the particle density, are reported in b), c) and e), f) for the patterns in a) and d), respectively. The reconstructed density of the xenon doping follows the color bar in Fig. 4. The shape of the helium droplets, independently retrieved before the imaging process and constrained using the DCDI method, are superimposed in blue color.
The experiment was performed at the European XFEL facility, taking advantage of the recently available capability of producing two collinear X-ray pulses of different radiation wavelengths with controllable time delay [19]. The two co-propagating pulses are tuned to $1.0\text{\,}\mathrm{keV}$ and $1.2\text{\,}\mathrm{keV}$ photon energy, corresponding to $1.2\text{\,}\mathrm{nm}$ and $1.0\text{\,}\mathrm{nm}$ wavelength, and can be separated in time from $50\text{\,}\mathrm{fs}$ up to $750\text{\,}\mathrm{fs}$. Both are focused in the interaction region, where they intercept isolated superfluid helium nanodroplets doped with xenon. An individual droplet in the focus interacts with both pulses, and the corresponding scattered fields incoherently overlap on the detector. Further details on the experiment are discussed in Sec. V.4.2.
Two diffraction images acquired in these conditions are shown in Fig. 2a and Fig. 2d for two different time delays between the two colors, i.e., $50\text{\,}\mathrm{fs}$ and $750\text{\,}\mathrm{fs}$, respectively. A relevant feature of these patterns is the low number of photons recorded by the detector, if compared to the photon yield for the standard, single pulse, operation mode of the XFEL [38] (see Sec. V.4.2). This aspect represents a challenge for a photon-hungry technique such as CDI, and thus even more so for Dichography (as further investigated in Sec. S1 of the Supplemental Material  [39]).
Nevertheless, the specific properties of the physical system under study, i.e., doped superfluid helium nanodroplets, allow us to take advantage of the Droplet Coherent Diffraction Imaging (DCDI) approach [40]. The DCDI method constrains the underlying helium droplet structure during the actual reconstruction process. The droplet size and brightness are extracted in the pre-processing phase, by fitting the experimental data with the diffraction profile of a spherical sample. The information to be retrieved, therefore, pertains solely to the internal structure of the doping material. For conventional CDI, this greatly reduces the complexity of the problem, it improves the convergence properties of the algorithms compared to standard CDI, and it makes the reconstruction process much more resilient to noise [40].
The DCDI method has been incorporated into the Dichography reconstruction process through a straightforward adaptation of its original workflow, as discussed in Sec. V.5, and is used here to analyze the two-color diffraction images presented in this section. Its applicability to this dataset relies on the possibility of extracting the droplet size as well as on the persistence of the pristine helium droplet density, embedding the xenon doping, up to a time delay of $750\text{\,}\mathrm{fs}$. Both aspects are investigated and verified using pure-helium data acquired during the same experimental campaign, as reported in Ref. [38].
Despite the demanding experimental conditions, the reduction in problem complexity achieved through the use of DCDI makes these two-color data tractable with Dichography in a few optimal cases.
The two frames reconstructed from the diffraction pattern shown in Fig. 2a are presented in Fig. 2b for the $1.0\text{\,}\mathrm{keV}$ pulse, and in Fig. 2c for the following $1.2\text{\,}\mathrm{keV}$ pulse. The two reconstructions are thus two snapshots of the same nanoparticle delayed by $50\text{\,}\mathrm{fs}$. In a similar way, the two frames reconstructed from the pattern in Fig. 2d correspond to the snapshots of the initial particle density (Fig. 2e) and its density after $750\text{\,}\mathrm{fs}$ (Fig. 2f).
In the reconstructions reported in Fig.  2, high density regions arise, structurally compatible with previous observations of xenon structures inside superfluid helium nanodroplets [41, 40, 42]. The first frame in Fig. 2b reveals the presence of four xenon agglomerates. The same structures can be identified in the second frame in Fig. 2b. However, the $1.2\text{\,}\mathrm{keV}$ reconstruction is characterized by a visibly lower quality, with “fuzzier” xenon agglomerates and the presence of a high density region on the left side of the droplet, which is not present in the $1.0\text{\,}\mathrm{keV}$ reconstruction. The lower quality of the $1.2\text{\,}\mathrm{keV}$ reconstruction can be explained with the imbalance of the scattering contributions: indeed, the recorded intensity of the scattered $1.2\text{\,}\mathrm{keV}$ signal is less than half of the $1.0\text{\,}\mathrm{keV}$ contribution.
[…]
From the representation of the retrieved densities in Fig. 2, the reader may question whether the similar structures in the two frames arise merely as artifacts, i.e., as a reproduction of the same signal in both frames. Here, we address this possibility as ghosting. We further discuss this aspect in the following Sec. II.2 and we provide a mathematical interpretation in Sec. V.2.4. In the case of two-color diffraction, i.e., of two different radiation wavelengths, the sample is imaged with intrinsically different spatial resolutions in the two frames. Specifically, the spatial extent corresponding to a single pixel varies between the frames, becoming smaller at higher photon energies due to the increased resolution at a shorter wavelength. As a result, the same structure spans a different number of pixels in the two reconstructions, and appears visually similar in size only when both are displayed with a common scale bar, as done in Fig. 2. This point is further illustrated in Fig. 7 and discussed in more detail in Sec. V.5. Ghosting in two-color Dichography would therefore give rise to a replicated structure but with a different scaling. Thus, its presence can be excluded for the xenon structures observed in Fig. 2.
The presence of the same structure in both frames, apart from the artifacts discussed above, implies that the xenon agglomerates survive for at least $750\text{\,}\mathrm{fs}$ after irradiation with the pump pulse. Thus, these results suggest that, at the irradiation level reached in this experiment, significant structural damage of the xenon filaments only happens at longer timescales. This observation, which holds up to the achieved resolution, is supported by molecular dynamics simulations [38].
While the geometry of the experiment would allow for a spatial resolution down to $5\text{\,}\mathrm{nm}$, the effective resolution achieved by the reconstructions is only around $20\text{\,}\mathrm{nm}$ (see Fig. S11 of the Supplemental Material). This is mostly due to the low intensity of the patterns that limits the maximum momentum transfer at which useful signal is recorded.
Among all the patterns recorded during the experiment, only a few turned out to have the right conditions for a dichographic reconstruction, due to the challenging conditions of the experiment. First of all, the pulse energy of the two-color pulses is reduced due to the specific XFEL configuration (see Sec. V.4.2). As a consequence, the brightness of the diffraction patterns is limited. In this low intensity condition, Dichography can successfully reconstruct the two frames only if their intensity contributions to the scattering signal are similar, as shown in Sec. S1.3 of the Supplemental Material [39]. A similar diffraction contribution requires a similar intensity of the two pulses when they interact with the droplet. This condition is however only achieved in a limited region of the focal volume due to the geometry of the experiment and the focusing optics, as also discussed in Sec. V.4.2. As a result, for this experiment, two-color diffraction patterns suitable for Dichography are “rare events” in the dataset and can be identified with a specific procedure described in Sec. S2 of the Supplemental Material [39].
II.2 Seeing double from a single shot
@@
Figure 3: Example of dichographic reconstruction from a single diffraction pattern produced by a “double-hit”. The experimental diffraction data from an individual light pulse delivered by SwissFEL is shown in a). The two particle densities retrieved from a) via Dichography are shown in b) and c). The two reconstructions correspond to two silver nanocubes, with edge length of around $100\text{\,}\mathrm{nm}$. Once the two particles have been reconstructed, the corresponding single-particle diffraction patterns can be calculated, shown in d) for the reconstruction in b), and in e) for the one reported in c). The central parts of the disentangled diffraction images in d) and e) are magnified in f) and g), respectively. The white arrows point at the same region of the diffraction image. All diffraction data in a), d), e), f) and g) are plotted with logarithmic color scale.
@@
The second set of results concerns the analysis of diffraction patterns produced by two particles struck by the same XFEL shot, known as “double-hits”, and they serve as a valuable probe for investigating the capabilities, performance, and key features of Dichography.
@@
Here, pairs of silver nanoparticles, simultaneously present in the XFEL focus, are intercepted by a single light pulse delivered by SwissFEL [43]. When the particles are far enough apart, the interference term between the two scattered fields is not resolved by the detector. The recorded diffraction image is, in practice, the incoherent sum of the two intensities, suitable for imaging with Dichography. The two retrieved frames correspond to the electron density of the two different particles. Further experimental information is provided in Sec. V.4.1.
@@
A first double-hit reconstruction is shown in Fig. 3. The images of the two particles in Fig. 3b and 3c are the two frames retrieved from the experimental diffraction pattern reported in Fig. 3a, without prior information regarding the particles’ shape, size, or orientation. The reconstructions reveal that the experimental image is produced by two silver nanocubes of slightly different sizes and alignment. Due to the cubic structure of the two samples and their orientation, each cube produces strong streaks with four-fold symmetry in the diffraction image. A total of eight streaks can be indeed identified in the diffraction pattern in Fig. 3a.
@@
It is possible to disentangle the two contributions to the scattering, i.e., to calculate the corresponding single-particle diffraction images from the two reconstructed frames. Fig. 3d and Fig. 3e report the two disentangled scattering contributions from the first (Fig. 3b) and second (Fig. 3c) frames, respectively. There, we can appreciate the capability of Dichography to separate the two contributions to the scattering superimposed in 3a, by reproducing the four-fold symmetry of the two cubes with different orientations. This situation is further muddled by a closer view in the central part of the separated patterns where the strongest contribution to the scattering signal is concentrated, provided in Fig. 3f and 3g.
@@
From these zoomed-in representations of the disentangled scattering signals, it is also possible to identify few regions where the intensities are not fully disentangled. For example, the white arrows in Fig. 3f and Fig. 3g indicate the same coordinates in the two patterns. In Fig. 3g, this area corresponds to a high intensity streak in the direction of cube facets, as visible from the reconstruction in Fig. 3c. Despite the absence of any related structural feature in the other reconstructed frame in Fig. 3b, its disentangled pattern in Fig. 3f still presents traces of this strong scattering signal. This signal marginally affects the inner density profile in Fig. 3c, slightly distorted in the direction of the cube facets of the other frame. We refer to this “cross-talking” between the two frames – where features in one create slight artifacts in the other due to strongly differing scattering intensity in specific directions – as ghosting. The origin of ghosting artifacts is further discussed in Sec. V.2.4. On the other hand, the intensities mistakenly assigned to the first frame in Fig. 3f are about two orders of magnitude lower than those correctly retrieved in Fig. 3g. This is evident from the logarithmic color map on the right and underscores the ability of Dichography to separate the two signals, even when their intensities locally differ by orders of magnitude in the diffraction signal.
@@
Further reconstruction attempts on diffraction patterns extracted from the same dataset of Fig. 3 are presented in Fig. 4. These examples are selected to highlight how Dichography performs on different sample configurations. For instance, Fig. 4a shows a reconstruction which reveals a pair of nanostructures with completely different geometries, a cube and a triangle, strongly underlining how this method neither requires any specific constraint in terms of sample’s structure, nor structural links between the two frames. The reconstruction in Fig. 4b has an additional challenging aspect. Here, the spatial extension of the two samples differs significantly, with the smaller one being around $100\text{\,}\mathrm{nm}$ in maximum size and the second reaching around $200\text{\,}\mathrm{nm}$. As a consequence, the smaller sample could spatially “fit” inside the larger one, potentially rendering the separation of the two frames of the reconstruction a significantly harder task. In both Fig. 4a and Fig. 4b, ghosting artifacts are barely visible in the form of line-shaped density fluctuations aligned with the hard edges of the reconstruction in the other frame.